Problem 3. If the HbS allele frequency is 0.105, compute the 
HbA/HbA, HbA/HbS, and HbS/HbS genotype frequencies.

Problem 4. If the HbA allele frequency is 0.196, compute the 
HbA/HbA, HbA/HbS, and HbS/HbS genotype frequencies.

Problem 5. If the HbS allele frequency is 0.158, compute the 
HbA/HbA, HbA/HbS, and HbS/HbS genotype frequencies.
Answer #1

3.

HbS = recessive allele = 0.105

HbA = dominant allele = 1 - 0.105 = 0.895

GF of HbA/HbA = 0.895 × 0.895 = 0.8010

GF of HbS/HbS = 0.105 × 0.105 = 0.111

GF of HbS/HbA = 2 × 0.105 × 0.895 = 0.1879

4.

HbA = 0.196

HbS = 1 - 0.196 = 0.804

GF of HbA/HbA = 0.196 × 0.196 = 0.0384

GF of HbS/HbS = 0.804 × 0.804 = 0.6464

GF of HbA/HbS = 2 × 0.196 × 0.804 = 0.3151

Please rate.